# Hui Xiong2, Bin Yang1

### v1

@@ Abstract
Anomaly detection in multivariate time series is challenging as heterogeneous subsequence anomalies may occur. Reconstruction-based methods, which focus on learning nomral patterns in the frequency domain to detect diverse abnormal subsequences, achieve promising resutls, while still falling short on capturing fine-grained frequency characteristics and channel correlations. To contend with the limitations, we introduce CATCH, a framework based on frequency patching. We propose to patchify the frequency domain into frequency bands, which enhances its ability to capture fine-grained frequency characteristics. To perceive appropriate channel correlations, we propose a Channel Fusion Module (CFM), which features a patch-wise mask generator and a masked-attention mechanism. Driven by a bi-level multi-objective optimization algorithm, the CFM is encouraged to iteratively discover appropriate patch-wise channel correlations, and to cluster relevant channels while isolating adverse effects from irrelevant channels. Extensive experiments on 9 real-world datasets and 12 synthetic datasets demonstrate that CATCH achieves state-of-the-art performance. We make our code and datasets available at https://anonymous.4open.science/r/CATCH-E535.
@@ 1 Introduction
Modern cyber physical systems are often monitored by multiple sensors, which produce sucessive multivariate time series data. Multivariate Time Series Anomaly Detection (MTSAD) aims to detect abnormal data in multivariate time series. It is applied widely including but not limited to financial fraud detection, medical disease identification, and cybersecurity threat detection (Li et al., 2021; Wen et al., 2022; Yang et al., 2023a; Kieu et al., 2018; 2019).
Time series anomalies are typically classified into point anomalies and subsequence anomalies. Recent reconstruction-based methods show strong capability of detecting point anomalies, which are characterized by specific values that significantly deviate from the normal range of the probability distribution. However, subsequence anomalies consist of values that fall within the probability distribution, making them much harder to detect (Paparrizos et al., 2022b; Nam et al., 2024). According to the behavior-driven taxonomy (Lai et al., 2021), the subsequence anomalies can be further divided into seasonal, shapelet, and trend anomalies–see Figure 1(a). A promising approach is to transform the time series into the frequency domain to better derive the subsequence anomalies.
@@
Figure 1: The frequency domain is partitioned into five frequency bands with the high-frequency bands zoomed in for clarity. (a) Shows a normal time series and its variations after being injected seasonal, shapelet, trend and mixture subsequence anomalies. In the time domain, anomalies are highlighted in red, while the affected frequency bands in the frequency domain are emphasized in gray. (b) Shows the frequency bands of a multivariate time series with three channels. These channels exhibit varying correlations across different frequency bands, each highlighted with distinct colors.
Figure 1(b) shows a multivariate time series with three channels, and we observe the varying channel associations in different frequency bands, where Channel 1 and Channel 2 are similar in the third band but dissimilar to Channel 3, and all channels are similar in the fourth band but show dissimilarity in the fifth. However, the commonly-used Channel-Independent (CI) and Channel-Dependent (CD) strategies exhibit polarization effects, rendering them inadequate for this task. CI uses the same model across different channels and overlooks potential channel correlations, which offers robustness (Nie et al., 2022) but lacks generalizability and capacity. CD considers all channels simultaneously with larger capacity, but may be susceptible to noise from irrelevant channels, thus lacking robustness (Han et al., 2024). This calls for flexibly adapting the distinct channel interrelationships in different frequency bands.
@@
  • We design the CFM to fully utilize the fine-grained channel correlations. Driven by a bi-level multi-objective optimization algorithm, the CFM is able to iteratively discover appropriate channel correlations and facilitate the isolation of irrelevent channels and the clustering of relevent channels, which provides both the capacity and robustness.
  • We conduct extensive experiments on 21 multivariate datasets. The results show that CATCH outperforms state-of-the-art baselines. Additionally, all datasets and code are avaliable at https://anonymous.4open.science/r/CATCH-E535.
@@ 2 Related Work
2.1 Multivariate Time-Series Anomaly Detection (MTSAD)
Traditional MTSAD methods can be classified into non-learning (Breunig et al., 2000; Goldstein & Dengel, 2012; Yeh et al., 2016) and machine learning (Liu et al., 2008; Ramaswamy et al., 2000). Recently, deep learning methods have shown exceptional MTSAD performance and have received substantial extensive attention. They can be classified into forecasting-based, reconstruction-based and contrastive-based methods. GDN (Deng & Hooi, 2021) is a forecasting-based model that uses a graph structure to learn topology and a graph attention network to encode input series, with anomaly detection based on the maximum forecast error among channel variables. Anomaly Transformer is a reconstructive approach that combines series and prior association to make anomalies distinctive (Xu et al., 2021). DCdetector uses contrastive learning in anomaly detection to create an embedding space where normal data samples are close together and anomalies are farther apart (Yang et al., 2023b). We focus on the reconstruction-based methods due to their prominent performance on commonly-used benchmark datasets (Qiu et al., 2024).
@@ 2.2 Channel Strategies in MTSAD
Regardless the above mentioned learning mechanism, a core issue is to model channel correaltions for better MTSAD. There are mainly two existing approaches that consider relationships among channels. Channel-Independent (CI) based methods such as: PatchTST (Nie et al., 2022) and DLinear (Zeng et al., 2023) impose the constraint of using the same model across different channels. While it offers robustness, it overlooks potential interactions among channels and can be limited in generalizability and capacity for unseen channels (Han et al., 2024). Previous studies have shown that correlation discovery in data is crucial for time series anomaly detection (Song et al., 2018). Channel-Dependent (CD) based method such as: MSCRED (Zhang et al., 2019) uses a convolutional-LSTM network with attention and a loss function to reconstruct correlation matrices among channels in multivariate time series input. iTransformer (Liu et al., 2023) embeds time points into variate tokens and applies an attention mechanism to capture multivariate correlations. MTAD-GAT (Zhao et al., 2020) treats each univariate time series as a feature and uses two parallel graph attention layers to capture dependencies across both temporal and channel dimensions. The existing methods could not adequately extract interrelationships, they may be susceptible to noise from irrelevant channels, reducing the model’s robustness.
2.3 Frequency Domain Analysis for MTSAD
Frequency-based models for MTSAD have gained significant attention in recent years. The Spectral Residual (SR) technique introduced in SR-CNN (Ren et al., 2019) leverages a frequency-based approach to create a saliency map for MTSAD. PFT (Park et al., 2021) uses partial Fourier transform to achieve substantial speedup without sacrificing accuracy. TFAD (Zhang et al., 2022) leverages frequency domain analysis with augmentation and decomposition. Nevertheless, the existing methods do not offer the point-granularity alignment for the frequency domain, still facing challenges in precisely detecting subsequence anomalies. Dual-TF (Nam et al., 2024) employs both the time and frequency domains for MTSAD, effectively addressing the granularity discrepancy between them. However, rather than introducing novel architectural changes, it adopts the existing anomaly-transformer (Xu et al., 2021) model. Our proposed CATCH offers fine-grained insights in the frequency domain and discovers appropriate channel correlations to fully reconstruct the normal patterns, thus strengthening the detection of subsequence anomalies.
@@ 3 CATCH
In the context of time series anomaly detection, the time series $\mathrm{X}\in\mathbb{R}^{N\times T}$ from the observed successive systems is represented by a sequence of equally spaced time points $<\mathrm{X}_{:,1},\mathrm{X}_{:,2},\cdots,\mathrm{X}_{:,T}>$, where $\mathrm{X}_{:,t}\in\mathbb{R}^{N},t=1,2,\cdots,T$ denotes the values of the observed $N$ features in time step $t$. The multivariate time series anomaly detection problem is to determine whether $\mathrm{X}_{:,t}$ is anomaly or not. For clear delineation, we separate dimensions with commas and use this format throughout this paper. For example, we denote $\mathrm{X}_{i,j}$ as the $i$-th channel at the $j$-th timestamp, $\mathrm{X}_{n,:}\in\mathbb{R}^{T}$ as the time series of $n$-th channel, where $n=1,2,\cdots,N$.
Figure 2: CATCH architecture. (1) Forward Module normalizes the input data, patchifies the frequency domain, and then projects it into the hidden space. (2) Channel Fusion Module captures channel interrelationships in each frequency band with a Channel-Masked Transformer (CMT) Layer, where the mask matrix (channel correlation) is generated by Mask Generator (MG). During backpropagation, MG is optimized by Channel Correlation Discovering mechanism for more appropriate channel correlations. (3) Time-Frequency Reconstruction Module obtains the frequency reconstruction through Flatten & Linear Head, and obtains the time reconstruction after iFFT.
@@ 3.1 Structure Overview
Figure 2 shows the overall architecture of the CATCH, which consists of three main moudles, the Forward Module, the Channel Fusion Module and the Time-Frequency Reconstruction Module.
In the Forward Module, we first apply the Instance Normalization Ulyanov et al. (2016) to unify the distribution of training and testing data. To model time series in both time and frequency domains, we then utilize the efficient FFT (Brigham & Morrow, 1967) to transform time series into orthogonal trigonometric signals in the frequency domain, where we keep both the real and imaginary (imag) parts through $\mathrm{X}^{R},\mathrm{X}^{I}=\text{FFT}(\mathrm{X})$ for maximum information retention, where $\mathrm{X},\mathrm{X}^{R},\mathrm{X}^{I}\in\mathbb{R}^{N\times T}$. Additionally, to capture fine-grained details in different frequency bands, we apply the patching operation in the frequency domain, the process is formalized as follows:
@@
where $\mathrm{P}^{R_{i}},\mathrm{P}^{I_{i}}\in\mathbb{R}^{N\times p}$ denote the $i$-th patch of $\mathrm{X}^{R}$ and $\mathrm{X}^{I}$. $L=[T-p]/s+1$ is the total patch number, where $p$ is the patch size and $s$ is the patch stride. We then concat each pair of $\mathrm{P}^{R_{i}}$ and $\mathrm{P}^{I_{i}}$ into $\mathrm{P}^{i}\in\mathbb{R}^{N\times 2p}$, as the $i$-th frequency patch. After patching in the frequency domain, the frequency patches are then projected into the high-dimensional hidden space through $\mathrm{P^{\prime}}^{i}=\text{Projection}(\mathrm{P}^{i})$.
Then, we utilize the Channel Fusion Module (CFM) to model the patch-wise channel correlations:
@@
where $\mathrm{P^{\prime}}^{i},\tilde{\mathrm{P}}^{i}\in\mathbb{R}^{N\times d}$. $d$ is the hidden dimension in attention blocks. The CFM parallels patch-wise to model the frequency patches simultaneously. We further introduce the details of CFM in Section 3.2 and the corresponding optimization mechanism in Section 3.3, which ensures the capability of CFM to discover appropriate channel correlations for fine-grained frequency bands.
Finally, we utilize the Time-Frequency Reconstruction Module (TFRM) to flatten the patch-wise representations and reconstruct all frequency spectrums with MLP projections separately for real and imaginary patches:
where $\mathrm{X^{\prime}}^{R},\mathrm{X^{\prime}}^{I}\in\mathbb{R}^{N\times T}$. And we also obtain the reconstruction in the time domain through $\mathrm{X^{\prime}}=\text{iFFT}(\mathrm{X^{\prime}}^{R},\mathrm{X^{\prime}}^{I})\in\mathbb{R}^{N\times T}$.
@@ 3.2 Channel Fusion Module
To better discriminate the heterogeneous subsequence anomalies with different occupancies on frequency bands, we devise the Channel Fusion Module which dynamically perceives the channel associations in each fine-grained frequency band and generates appropriate channel correlations to enhance the reconstruction ability. Specifically, we propose a patch-wise mask generator and utilize the masked attention mechanism to capture the channel correlations.
Patch-wise Mask Generator. Inspired by Selective State Space Models such as Mamba (Gu & Dao, 2023; Dao & Gu, 2024), which utilizes Linear projections to flexibly update the hidden states based on the current data for larger capacity, the patch-wise channel associations can also be seen as a changing hidden state strongly associated with the current patch. Therefore, we devise a Linear-based mask generator to perceive the suitable channel associations for each frequency band by generating binary mask matrices to isolate the adverse effects from irrelevant channels. Note that the binary mask is an intermediate state between CI (identity matrix) and CD (all-ones matrix) strategies. Moreover, the mask generator itself works in a CI manner to mitigate the adverse effects from noisy channels. To ensure the rationality of discovered channel correlations and provide both robustness and capacity, we also design specific optimization objectives in Section 3.3. We take the $i$-th frequency patch as an example:
@@
where $\mathcal{D}^{i}$ is the probability matrix for $i$-th patch, with the shape of $\mathbb{R}^{N\times N}$. Sigmoid function $\sigma$ projects the values to probabilities. Since our goal is to filter out the adverse effects of irrelevant channels, we further perform Bernoulli resampling on the probability matrices to obtain binary mask matrix $\mathcal{M}^{i}$ with the same shape. Higher probability $\mathcal{D}^{i}_{l,m}$ results in $\mathcal{M}^{i}_{l,m}$ closer to 1, indicating a relationship between channel $l$ and channel $m$. And we manually keep the diagonal items to $1$. To ensure the propagation of gradients, we use the Gumbel Softmax reparameterization trick (Jang et al., 2016) during Bernoulli resampling.
Channel-Masked Transformer Layer. After the patch-wise mask generator outputs the mask matrices for frequency bands, we utilize the transformer layer to further capture the fine-grained channel correlations. The Layer Normalization is applied before each attention block to mitigate the over-focusing phenomenon on frequency components with larger amplitudes (Piao et al., 2024):
$\mathrm{P^{\ast}}^{i}=\text{LayerNorm}(\mathrm{P^{\prime}}^{i})=(\mathrm{P^{\prime}}^{i}-\text{Mean}_{n=1}^{N}(\mathrm{P^{\prime}}_{n,:}^{i}))/\sqrt{\text{Var}_{n=1}^{N}(\mathrm{P^{\prime}}_{n,:}^{i})}$ (6)
Empirically, we utilize the masked attention mechanism to further model the fine-grained interrelationships among relevent channels and integrate the mask in a calculated way to keep the propagation of gradients:
@@
where $\mathrm{P^{\ast}}^{i}\in\mathbb{R}^{N\times d}$, $\mathrm{\tilde{P}}^{i}\in\mathbb{R}^{N\times d}$, $\mathrm{W}^{Q},\mathrm{W}^{K},\mathrm{W}^{V}\in\mathbb{R}^{d\times d}$. We utilize the same Feed-Forward networks and skip connections as the classical transformers (Vaswani, 2017). We also apply multi-head mechanism to jointly attend to information from different representational subspaces, and the Channel-Masked Transformer Layer can be stacked multiple times.
3.3 Time-Frequency Training
We design a novel training process to enhance the model’s ability to detect both point anomalies and subsequence anomalies. And the Channel Fusion Module (CFM) is also encouraged to discover appropriate channel correlations for different frequency bands. Specifically, we design multiple optimization objectives and a bi-level optimization process to achieve these.
@@ 3.3 Time-Frequency Training
Fine-grained Channel Correlation Discovering. We introduce the forward propagation process of CFM in Section 3.2. From an optimization perspective, it is essential to design appropriate optimization objectives to enhance the effectiveness of generated masks. A direct motivation is to explicitly enhance the attention scores between relevent channels defined by the mask, thus aligning the attention mechanism with the currently discovered optimal channel correlation, which helps isloate the adverse effects from irrelevant channels and provides robustness for the attention mechanism. Then we iteratively optimize the mask generator to refine the channel correlations, tuning the capacity of attention mechanism to fully capture the interrelationships between channels. Intuitively, we devise a compound loss function to guide the mask generator exploring the space of channel correlations (from CI to CD). The proposed compound loss function for patch ${\mathrm{P}^{\ast}}^{i}$ is formalized as:
$\text{ClusteringLoss}=-\frac{1}{N}\displaystyle\sum_{k=1}^{N}\log\bigg{(}\frac{\sum_{m=1}^{N}exp(\mathcal{S}^{i}_{k,m})/\tau}{\sum_{l=1}^{N}exp(\mathcal{T}^{i}_{k,l})/\tau}\bigg{)},\text{RegularLoss}=\frac{1}{N}||\mathrm{I}-\mathcal{M}^{i}||_{F},$ (10)

### v3

@@ Abstract
Anomaly detection in multivariate time series is challenging as heterogeneous subsequence anomalies may occur. Reconstruction-based methods, which focus on learning normal patterns in the frequency domain to detect diverse abnormal subsequences, achieve promising results, while still falling short on capturing fine-grained frequency characteristics and channel correlations. To contend with the limitations, we introduce CATCH, a framework based on frequency patching. We propose to patchify the frequency domain into frequency bands, which enhances its ability to capture fine-grained frequency characteristics. To perceive appropriate channel correlations, we propose a Channel Fusion Module (CFM), which features a patch-wise mask generator and a masked-attention mechanism. Driven by a bi-level multi-objective optimization algorithm, the CFM is encouraged to iteratively discover appropriate patch-wise channel correlations, and to cluster relevant channels while isolating adverse effects from irrelevant channels. Extensive experiments on 10 real-world datasets and 12 synthetic datasets demonstrate that CATCH achieves state-of-the-art performance. We make our code and datasets available at https://github.com/decisionintelligence/CATCH.
@@ 1 Introduction
Modern cyber physical systems are often monitored by multiple sensors, which produce sucessive multivariate time series data (Wang et al., 2024e; Dai et al., 2024; Sun et al., 2025; Tian et al., 2025; Guo et al., 2014). In recent years, there has been significant progress in multivariate time series analysis, with key tasks such as forecasting (Qiu et al., 2025c; Yu et al., 2023; Wu et al., 2024a; Liu et al., 2024a; Huang et al., 2024; Chen et al., 2024), classification (Yao et al., 2024b; Campos et al., 2023), and imputation (Gao et al., 2024; Wang et al., 2024c; b), among others (Huang et al., 2023; Yao et al., 2024a; Liu et al., 2025), gaining attention. Among these, Multivariate Time Series Anomaly Detection (MTSAD), which focuses on identifying abnormal data in multivariate time series, stands out as a critical studied task. It is applied widely including but not limited to financial fraud detection, medical disease identification, and cybersecurity threat detection (Wen et al., 2022; Yang et al., 2023a; Kieu et al., 2018; 2019; Wu et al., 2024b; Shentu et al., 2024; Yang et al., 2021).
Time series anomalies are typically classified into point anomalies and subsequence anomalies. The point anomalies can be further classified as contextual or global anomalies (Lai et al., 2021). Recent reconstruction-based methods show strong capability of detecting point anomalies, which are characterized by specific values that significantly deviate from the normal range of the probability distribution. However, subsequence anomalies consist of values that fall within the probability distribution, making them much harder to detect (Paparrizos et al., 2022b; Nam et al., 2024). According to the behavior-driven taxonomy (Lai et al., 2021), the subsequence anomalies can be further divided into seasonal, shapelet, and trend anomalies–see Figure 1(a). A promising approach is to transform the time series into the frequency domain to better derive the subsequence anomalies.
@@
Figure 1: The frequency domain is partitioned into five frequency bands with the high-frequency bands zoomed in for clarity. (a) Shows a normal time series and its variations after being injected seasonal, shapelet, trend and mixture subsequence anomalies. In the time domain, anomalies are highlighted in red, while the affected frequency bands in the frequency domain are emphasized in gray. (b) Shows the frequency bands of a multivariate time series with three channels. These channels exhibit varying correlations across different frequency bands, where Channels 1 and 2 exhibit similar behavior in the third frequency band and are therefore marked in red, while Channel 3 exhibits distinct characteristics and is marked in yellow. In the fourth frequency band, all channels behave similarly and are marked in green, while in the fifth band, they exhibit distinct characteristics and are marked in different colors.
Figure 1(b) shows a multivariate time series with three channels, and we observe the varying channel associations in different frequency bands, where Channel 1 and Channel 2 are similar in the third band but dissimilar to Channel 3, and all channels are similar in the fourth band but show dissimilarity in the fifth. However, the commonly-used Channel-Independent (CI) and Channel-Dependent (CD) strategies exhibit polarization effects, rendering them inadequate for this task. CI uses the same model across different channels and overlooks potential channel correlations, which offers robustness (Nie et al., 2023; Qiu et al., 2025a) but lacks generalizability and capacity. CD considers all channels simultaneously with larger capacity, but may be susceptible to noise from irrelevant channels, thus lacking robustness (Han et al., 2024; Qiu et al., 2025a). This calls for flexibly adapting the distinct channel interrelationships in different frequency bands.
@@
  • We design the CFM to fully utilize the fine-grained channel correlations. Driven by a bi-level multi-objective optimization algorithm, the CFM is able to iteratively discover appropriate channel correlations and facilitate the isolation of irrelevant channels and the clustering of relevant channels, which provides both the capacity and robustness.
  • We conduct extensive experiments on 22 multivariate datasets. The results show that CATCH outperforms state-of-the-art baselines.
@@ 2 Related Work
2.1 Multivariate Time-Series Anomaly Detection
Traditional MTSAD methods can be classified into non-learning (Breunig et al., 2000; Goldstein & Dengel, 2012; Yeh et al., 2016) and machine learning (Liu et al., 2008; Ramaswamy et al., 2000). Recently, deep learning methods have shown exceptional MTSAD performance and have received substantial extensive attention (Liu et al., 2024b; Liu & Paparrizos, 2024; Hu et al., 2024; Wang et al., 2023; Li et al., 2021; Wang et al., 2024a; f). They can be classified into forecasting-based, reconstruction-based and contrastive-based methods. GDN (Deng & Hooi, 2021) is a forecasting-based model that uses a graph structure to learn topology and a graph attention network to encode input series, with anomaly detection based on the maximum forecast error among channel variables. Anomaly Transformer is a reconstructive approach that combines series and prior association to make anomalies distinctive (Xu et al., 2021). DCdetector uses contrastive learning in anomaly detection to create an embedding space where normal data samples are close together and anomalies are farther apart (Yang et al., 2023b).
@@ 2.2 Channel Strategies in MTSAD
A channel refers to a variable in MTSAD, while a channel strategy refers to how these channel correlations are effectively considered during the modeling process. CATCH employs a reconstruction-based anomaly detection algorithm, using reconstruction error as the anomaly score, making reconstruction quality crucial for anomaly detection accuracy. Since the channels in multivariate time series often exhibit complex dependencies, explicitly modeling these correlations enables a more comprehensive capture of global features, thereby improving reconstruction capabilities and anomaly detection performance. There are mainly two existing approaches that consider relationships among channels. Channel-Independent (CI) based methods such as: PatchTST (Nie et al., 2023) and SparseTSF (Lin et al., 2024) impose the constraint of using the same model across different channels. While it offers robustness, it overlooks potential interactions among channels and can be limited in generalizability and capacity for unseen channels (Han et al., 2024). Previous studies have shown that correlation discovery in data is crucial for time series anomaly detection (Song et al., 2018). Channel-Dependent (CD) based methods such as: MSCRED (Zhang et al., 2019) uses a convolutional-LSTM network with attention and a loss function to reconstruct correlation matrices among channels in multivariate time series input. iTransformer (Liu et al., 2024c) embeds time points into variate tokens and applies an attention mechanism to capture multivariate correlations. MTAD-GAT (Zhao et al., 2020) treats each univariate time series as a feature and uses two parallel graph attention layers to capture dependencies across both temporal and channel dimensions. The existing methods could not adequately extract interrelationships, they may be susceptible to noise from irrelevant channels, reducing the model’s robustness.
2.3 Frequency Domain Analysis for TSAD
Table 1: Comparison of existing frequency-based TSAD methods.
Property Mmultivariate time serie anomaly detection Time-frequency granularity alignment Handle high-frequency information Capture channel correlations
SR-CNN
PFT
TFAD
Dual-TF
CATCH
Frequency domain analysis can uncover subsequence anomalies that are challenging to detect in the time domain, such as anomalies in periodic fluctuations or oscillation patterns, significantly enhancing detection accuracy (Zhang et al., 2022). Consequently, frequency-based time series anomaly detection models have garnered widespread attention in recent years. SR-CNN (Ren et al., 2019), the first method to leverage the frequency domain for TSAD, employs a frequency-based approach to generate saliency maps for identifying anomalies, and PFT (Zhang et al., 2022) built on this foundation by introducing partial Fourier transform to achieve substantial acceleration. However, both methods are confined to univariate time series and fail to address the complexities of multivariate scenarios. TFAD (Zhang et al., 2022) emerges as the first approach to integrate time-domain and frequency-domain analyses for MTSAD, yet it lacks time-frequency granularity alignment. Dual-TF (Nam et al., 2024), the most recent algorithm for MTSAD using time-frequency analysis, partially addresses the time-frequency granularity alignment issue but lacks a tailored backbone, instead relying directly on the Anomaly Transformer (Xu et al., 2021). While TFAD and Dual-TF represent progress in this area, they still exhibit the following limitations: i) Frequency-domain modeling methods have inherent biases, often overlooking high-frequency information; 2) Insufficient exploration and utilization of channel correlations in multivariate time series.
@@ 3 CATCH
In the context of time series anomaly detection, $\mathrm{X}\in\mathbb{R}^{N\times T}$ denotes a time series with $N$ channels and length $T$. For clear delineation, we separate dimensions with commas and use this format throughout this paper. For example, we denote $\mathrm{X}_{i,j}$ as the $i$-th channel at the $j$-th timestamp, $\mathrm{X}_{n,:}\in\mathbb{R}^{T}$ as the time series of $n$-th channel, where $n=1,2,\cdots,N$. The multivariate time series anomaly detection problem is to determine whether $\mathrm{X}_{:,t}$ is anomaly or not.
Figure 2: CATCH architecture. (1) Forward Module normalizes the input data, patchifies the frequency domain, and then projects it into the hidden space. (2) Channel Fusion Module captures channel interrelationships in each frequency band with a Channel-Masked Transformer (CMT) Layer, where the mask matrix (channel correlation) is generated by Mask Generator (MG). During training, MG and CMT are optimized by Channel Correlation Discovering (CCD) for more appropriate channel correlations. (3) Time-Frequency Reconstruction Module obtains the frequency reconstruction through Flatten & Linear Head Layer, and obtains the time reconstruction after iFFT.
@@ 3.1 Structure Overview
Figure 2 shows the overall architecture of the CATCH, which consists of three main modules: 1) the Forward Module, 2) the Channel Fusion Module (CFM), and 3) the Time-Frequency Reconstruction Module (TFRM). Specifically, the input multivariate time series is firstly processed via the Forward Module, consisting of the Instance Norm Layer, FFT&Patching Layer, and Projection Layer.
The Instance Norm Layer is firstly to mitigate the distributional shifts between the training and testing data caused by varying statistical properties, enhancing the model’s generalization in reconstructing the testing data. Then, the FFT&Patching Layer is to perform fine-grained modeling in each frequency band (patch). Specifically, we utilize the efficient FFT (Brigham & Morrow, 1967) to transform time series into orthogonal trigonometric signals in the frequency domain, where we keep both the real and imaginary parts through $\mathrm{X}^{R},\mathrm{X}^{I}=\text{FFT}(\mathrm{X})$ for maximum information retention, where $\mathrm{X},\mathrm{X}^{R},\mathrm{X}^{I}\in\mathbb{R}^{N\times T}$. Next, we apply the frequency patching operation to create fine-grained frequency bands (patches), which can be formalized as follows.
@@
where $\mathrm{P}^{R_{i}},\mathrm{P}^{I_{i}}\in\mathbb{R}^{N\times p}$ denote the $i$-th patch of $\mathrm{X}^{R}$ and $\mathrm{X}^{I}$. $L=[T-p]/s+1$ is the total patch number, where $p$ is the patch size and $s$ is the patch stride.
We then concat each pair of $\mathrm{P}^{R_{i}}$ and $\mathrm{P}^{I_{i}}$ into $\mathrm{P}^{i}\in\mathbb{R}^{N\times 2p}$, as the $i$-th frequency patch. After patching in the frequency domain, the frequency patches are then projected into the high-dimensional hidden space through the Projection Layer: $\mathrm{P^{\prime}}^{i}=\text{Projection}(\mathrm{P}^{i})$.
After the Forward Module, the processed time series is fed into the Channel Fusion Module (CFM) to dynamically model the channel correlations in each fine-grained frequency band.
@@
where $\mathrm{P^{\prime}}^{i},\tilde{\mathrm{P}}^{i}\in\mathbb{R}^{N\times d}$, $N$ is the number of channels and $d$ is the hidden dimension in attention blocks. The CFM parallels in a patch-wise way to model the frequency patches simultaneously. We further introduce the details of CFM in Section 3.2.
Next, we utilize the Time-Frequency Reconstruction Module (TFRM) to reconstruct all frequency spectrums for real and imaginary patches and simultaneously obtain their temporal reconstruction:
$\mathrm{X^{\prime}},\mathrm{X^{\prime}}^{R},\mathrm{X^{\prime}}^{I}=\text{TFRM}(\{\mathrm{\tilde{P}}^{R_{1}},\mathrm{\tilde{P}}^{R_{2}},\cdots,\mathrm{\tilde{P}}^{R_{L}}\},\{\mathrm{\tilde{P}}^{I_{1}},\mathrm{\tilde{P}}^{I_{2}},\cdots,\mathrm{\tilde{P}}^{I_{L}}\}),$ (3)
where $\mathrm{X^{\prime}}$ is the temporal reconstruction, and $\mathrm{X^{\prime}}^{R},\mathrm{X^{\prime}}^{I}\in\mathbb{R}^{N\times T}$ are the frequency reconstruction. Finally, we integrate the reconstruction error in both time and frequency domains as anomaly score.
@@ 3.2 Channel Fusion Module
The Channel Fusion Module (CFM) contains the following three components: 1) the Mask Generator (MG), 2) the Channel-Masked Transformer Layer (CMT), and 3) the Channel Correlation Discovering (CCD) mechanism. Specifically, the MG is to perceive and generate the mask matrices (channel correlations) in different frequency bands and guide the masked attention in CMT. The CMT aims to model the appropriate patch-wise channel correlations. And the CCD is to guide the MG and CMT to explore better channel correlations during optimization.
Mask Generator. Inspired by Selective State Space Models such as Mamba (Dao & Gu, 2024), which utilizes Linear projections to flexibly update the hidden states based on the current data for larger capacity, the patch-wise channel associations can also be treated as a changing hidden state strongly associated with the current patch. Therefore, we devise a Linear-based mask generator to perceive the suitable channel associations for each frequency band by generating binary mask matrices to isolate the adverse effects from irrelevant channels. Note that the binary mask is an intermediate state between CI (identity matrix) and CD (all-ones matrix) strategies. We take the $i$-th frequency patch as an example:
@@
where $\mathrm{P^{\prime}}^{i}\in\mathbb{R}^{N\times d}$, $\mathcal{D}^{i}\in\mathbb{R}^{N\times N}$, and $\mathcal{M}^{i}\in\mathbb{R}^{N\times N}$ are the hidden representation, the probability matrix, and the binary mask matrix of $i$-th patch, respectively. $\sigma$ projects the values to probabilities.
Since our goal is to filter out the adverse effects of irrelevant channels, we further perform Bernoulli resampling on the probability matrices to obtain binary mask matrix $\mathcal{M}^{i}$ with the same shape. Higher probability $\mathcal{D}^{i}_{l,m}$ results in $\mathcal{M}^{i}_{l,m}$ closer to 1, indicating a relationship between channel $l$ and channel $m$. And we manually keep the diagonal items to $1$. To ensure the propagation of gradients, we use the Gumbel Softmax reparameterization trick (Jang et al., 2017) during Bernoulli resampling.
Channel-Masked Transformer Layer. After the Mask Generator outputs the mask matrices for frequency bands, we utilize the transformer layer to further capture the fine-grained channel correlations. The Layer Normalization is applied before each attention block to mitigate the over-focusing phenomenon on frequency components with larger amplitudes (Piao et al., 2024):
$\mathrm{P^{\ast}}^{i}=\text{LayerNorm}(\mathrm{P^{\prime}}^{i})=(\mathrm{P^{\prime}}^{i}-\text{Mean}_{n=1}^{N}(\mathrm{P^{\prime}}_{n,:}^{i}))/\sqrt{\text{Var}_{n=1}^{N}(\mathrm{P^{\prime}}_{n,:}^{i})},$ (5)
where $\mathrm{P^{\prime}}^{i}\in\mathbb{R}^{N\times d}$ and $\mathrm{P^{\ast}}^{i}\in\mathbb{R}^{N\times d}$ are the hidden representation and the normalized representation of $i$-th patch, respectively. Empirically, we utilize the masked attention mechanism to further model the fine-grained interrelationships among relevant channels and integrate the mask matrix in a calculated way to keep the propagation of gradients:
@@
where $\mathrm{W}^{Q},\mathrm{W}^{K},\mathrm{W}^{V}\in\mathbb{R}^{d\times d}$. $\mathcal{M}^{i}\in\mathbb{R}^{N\times N}$, $\mathcal{T}^{i}\in\mathbb{R}^{N\times N}$, $\mathcal{S}^{i}\in\mathbb{R}^{N\times N}$, and $\mathrm{\tilde{P}}^{i}\in\mathbb{R}^{N\times d}$ are the binary mask matrix, the attention matrix, the masked attention matrix, and the hidden representation processed by CMT of $i$-th patch, respectively.
@@
We utilize the same Feed-Forward networks and skip connections as the classical transformers (Vaswani et al., 2017). We also apply multi-head mechanism to jointly attend to information from different representational subspaces and the CMT can be stacked multiple times.
@@
Channel Correlation Discovering. From an optimization perspective, it is essential to design appropriate optimization objectives to enhance the effectiveness of generated masks. A direct motivation is to explicitly enhance the attention scores between relevant channels defined by the mask, thus aligning the attention mechanism with the currently discovered optimal channel correlation, which helps isloate the adverse effects from irrelevant channels and provides robustness for the attention mechanism. Then we iteratively optimize the Mask Generator to refine the channel correlations, tuning the capacity of attention mechanism in the Channel-Masked Transformer Layer to fully capture the interrelationships between channels. Intuitively, we devise two loss functions to guide the Mask Generator exploring the space of channel correlations (from CI to CD). The proposed loss functions are formalized as: $\text{ClusteringLoss}=-\frac{1}{N}\displaystyle\sum_{k=1}^{N}\log\bigg{(}\frac{\sum_{m=1}^{N}exp(\mathcal{S}^{i}_{k,m})/\tau}{\sum_{l=1}^{N}exp(\mathcal{T}^{i}_{k,l})/\tau}\bigg{)},$ (9) $\text{RegularLoss}=\frac{1}{N}||\mathrm{I}-\mathcal{M}^{i}||_{F},$ (10)
where $\tau$ is the temperature coefficient, $N$ is the number of channels, $\mathrm{I}$ is the identity matrix, and $||\cdot||_{F}$ is the Frobenius norm. $\mathcal{T}^{i}\in\mathbb{R}^{N\times N}$, $\mathcal{S}^{i}\in\mathbb{R}^{N\times N}$, and $\mathcal{M}^{i}\in\mathbb{R}^{N\times N}$ are the attention matrix, the masked attention matrix, and the binary mask matrix for $i$-th patch, respectively.